# references.bib Weak Ties Explain Open Source Innovation

### v1

@@
Abstract
In a real-world social network, weak ties (reflecting low-intensity, infrequent interactions) act as bridges and connect people to different social circles, giving them access to diverse information and opportunities that are not available within one’s immediate, close-knit vicinity. Weak ties can be crucial for creativity and innovation, as it introduces new ideas and approaches that people can then combine in novel ways, leading to innovative solutions and creative breakthroughs. Do weak ties facilitate creativity in software in similar ways?
@@ Abstract
In this paper, we show that the answer is “yes.” Concretely, we study the correlation between developers’ knowledge acquisition through three distinct interaction networks on GitHub and the innovativeness of the projects they develop, across over 38,000 Python projects hosted on GitHub. Our findings suggest that the diversity of projects in which developers engage correlates positively with the innovativeness of their future project developments, whereas the volume of interactions exerts minimal influence. Notably, acquiring knowledge through weak interactions (e.g., starring) as opposed to strong ones (e.g., committing) emerges as a stronger predictor of future novelty.
I Introduction
In the late 1960s and early 1970s, Mark Granovetter, then a PhD student at Harvard University, studied a random sample of men living in a Boston suburb who had recently changed jobs, to understand how they learned about their new job opportunities. His 1973 paper [granovetter1973strength], which would go on to become the most cited work in the social sciences, had one key finding: that while a significant majority of people found their jobs through personal contacts rather than through formal channels like advertisements or employment agencies, these contacts were acquaintances (weak ties) rather than close friends or family (strong ties). Granovetter’s resulting theory on the “strength of weak ties” explains how weak ties are more effective in job searches—they act as bridges and connect individuals to different social circles, giving access to diverse information and opportunities that are not available within one’s immediate, close-knit network. The theory continues to accrue scholarly attention and large-scale confirmation in the digital age [rajkumar2022causal].
Beyond job searches, the theory had far-reaching implications. For example, weak ties impact, the diffusion of information, social mobility, community organization, and, the focus of this paper, creativity and innovation. The greater diversity of information enabled by weak ties can be crucial for creativity and innovation, as it introduces new ideas and approaches that might not be available within a close-knit group of strong ties, where information is more likely to be redundant. People can then combine these ideas in novel ways, leading to innovative solutions and creative breakthroughs [baer2010strength, rost2011strength, michelfelder2013and]. The famous “water-cooler effect,” that advocates of in-person work environments often quote, relies on the same mechanism—these casual encounters, which are typically with weak ties (work acquaintances rather than close colleagues), can lead to the exchange of diverse ideas and perspectives that spark creativity [brucks2022virtual].
Do weak ties facilitate creativity in software in similar ways? More generally, how are structural characteristics of the information networks of software creators associated with the innovativeness of the software they will create? For example, is the software more innovative when it is created by developers with access to more diverse information networks?
In this paper we answer these questions quantitatively, through a large-scale empirical study of over 38,000 open-source Python projects hosted on GitHub. For every project in our sample, we start by reconstructing three project-to-project networks based on interactions (i.e., commits, issues, and GitHub star events) of its core developers with other projects in the past, spanning ties of varying strength. Using these networks we then compute proxies for the amount and diversity of knowledge accessible to the core developers, before working on the focal project, via their weak and strong ties. Finally, building on a recent framework viewing software innovation as the novel recombination of software packages [fang2024novelty], we propose a novel measure of software innovativeness and estimate regression models to quantify the association between this measure and various network characteristics.
Our results indicate that software innovativeness increases when actors have access to a greater diversity of projects through the three networks we analyze, whereas the volume of projects they interacted with in the past does not correlate with project innovativeness. Moreover, we observe that the diversity of projects and knowledge accessed through low-effort interactions (weak ties), as captured for instance by “starring” repositories, has a greater impact on subsequent project innovativeness compared to interactions involving higher-effort engagements (strong ties) like issue comments and commits.
Overall, our study makes several key contributions, including a novel measure of software innovativeness, and an empirically validated operationalization of tie strength derived from actions available on GitHub. But most notably, we begin to fill an important knowledge gap. Despite growing realization of the importance of better understanding and enabling creativity and innovation in software engineering [jackson2022team, jackson2024creativity, inman2024developer], there is a dearth of research and empirical evidence in our community on the mechanisms through which creativity and innovation emerge. Our work is the first to translate established social science theory on the strength of weak ties to the software engineering domain, and to provide strong empirical evidence that similar innovation-enabling mechanisms apply also in (open-source) software. We discuss more implications for platform design and open source practice in Section V.
II Theoretical Framework and Hypotheses
In this section we develop our research hypotheses, drawing on established social science theories and a rich literature on open-source development and the GitHub platform.
II-A Innovation as Novel Recombination
Innovation is a critical aspect of science and technology development, marked by profound changes that can disrupt established competencies and lead to significant advancements [verhoeven2016measuring]. Open-source software is a manifestation of open innovation [west2006challenges, west2006patterns], by fostering a collaborative, transparent, and user-driven development environment that uses the collective intelligence and resources of a globally-distributed community. This approach accelerates the pace of innovation and leads to more robust, adaptable, and widely adopted software solutions that are increasingly being recognized as critical digital infrastructure that powers our society [eghbal2020working].
There have been various attempts to characterize innovation in software, with early work [couger1992measurement] recognizing the dimensions of novelty and utility and more recent work also considering distinctions in functionality, aesthetics, and other relevant features [beghelli2020novelty]. One prevalent discourse within the scholarly literature on scientific and technological innovation asserts that innovation often arises from preexisting components of knowledge [uzzi2013atypical, shi2023surprising], with the reconfiguration of knowledge serving as a predominant mechanism for generating novel and useful outcomes [schumpeter1939business]. In open source, this framing led to recently establishing an empirical link between the integration of atypical combinations of software packages, indicating novelty, and the higher subsequent popularity of the corresponding software projects, as a manifestation of utility [fang2024novelty].
How to facilitate the emergence of innovation remains unclear, though a common argument centers on facilitating creativity, as a precursor. Creativity and innovation are closely related but distinct concepts—creativity is about generating novel ideas, while innovation is about applying those ideas to create real value. It is expected that a software team that is more creative has access to richer knowledge (e.g., knowledge about the problem and technical approaches to solving it), and is better able to combine pieces of that knowledge to create an innovative solution [du2007role].
@@ II-A Innovation as Novel Recombination
Of course, there are many facets to creativity, including as an individual personality trait [kandler2016nature]. In this paper we focus only on one. Specifically, we argue that creativity at the team level is a function of the knowledge networks of the team members and, while we cannot measure creativity directly, we can expect to see differences in the teams’ creative outputs (i.e., more novel software), that are associated with differences in the structure of those knowledge networks.
II-B Knowledge Acquisition through Social Learning
@@
Historically, and especially with the advent of “social coding” platforms like GitHub, participation in open-source development has been rife with opportunities for social learning [crowston2005social], i.e., the process by which individuals acquire new knowledge, behaviors, skills, or attitudes through observation, imitation, and interaction with others. First, learning, including by observing what others are doing [yang2021differential], has been [lakhani2005hackers, krishnamurthy2006intrinsic] and remains [gerosa2021shifting] among the most important motivations for people to contribute to open source. In addition, modern code hosting platforms offer many opportunities for users to interact (e.g., collaboration on a shared codebase, issue discussions, code review. Through social media-like functionality (e.g., “following” [blincoe2016understanding]) and many available signals that offer a high degree of transparency [dabbish2012social] (e.g., repository badges [trockman2018adding]), the platforms also facilitate users quietly observing and being influenced by others’ behaviors.
@@
These interactions with other individuals [von2003community] and the engagement with the artifacts they create [cervetti2009increasing] serve as channels for information dissemination and knowledge exchange. Indeed, prior studies have documented many such social learning effects, including choosing an open-source license [peng2013learning, singh2013networks], discovering and adopting emerging tools [singer2014software, lamba2020heard], and learning new software design principles and programming skills by reviewing code authored by others [nandigam2008learning, alami2019does].
@@
[…]
@@ II-C The Strength of Weak Ties
That in a social network ties can have varying strength, reflective of real-world factors like the duration of shared interactions, emotional depth, level of intimacy, and amount of reciprocal exchanges [granovetter1973strength], should be no surprise. Consequently, one can thus label the ties involving infrequent and less intimate interactions as “weak,” and those characterized by higher frequency of interaction and intimacy as “strong.”
Where Granovetter’s theory [granovetter1973strength] was groundbreaking, however, is in explaining that weak ties play a fundamentally different structural role in networks. As it turns out, strong ties tend to develop among individuals within the same community, while weak ties are more common between individuals from different communities. Here, a community is defined as a group of people who are predominantly interconnected with each other. This pattern arises because individuals within the same community have numerous opportunities for interaction and often share many mutual friends. This shared social context leads to increased similarity in behaviors, interests, and other characteristics, thereby enhancing intimacy among individuals [granovetter1973strength]. Furthermore, Heider’s Cognitive Balance Theory [munroe2007cognitive] proposes that when two individuals are linked only by weak ties yet share many common friends, it creates a “psychological strain;” and such situations, characterized by cognitive imbalance, are typically avoided by people. Therefore, weak ties between individuals sharing many mutual friends or belonging to the same community are unlikely to persist over time [granovetter1973strength].
The bridging role of weak ties between people of different communities offers great advantages for the diffusion of novel information. Within the confines of a single community, individuals tend to become more similar through frequent social interactions, leading to the consumption and circulation of similar types of information. Consequently, the information disseminated with the strong ties connecting individuals of a single community often contains less novel content, and at times may even be redundant, rendering it less valuable overall (the discussion of echo chambers on social media [kitchens2020understanding] is a prime example of this degradation of information quality). In contrast, information acquired through weak ties, or cross-community connections, tends to originate more from individuals with diverse backgrounds and knowledge bases. This information is often unavailable within one’s own community, making it more likely to be novel and of greater value [granovetter1973strength].
@@
The advantage of weak ties in accessing novel and valuable information has been empirically evaluated in various contexts, including job seeking [rajkumar2022causal, granovetter1983strength], generation of innovative ideas [burt2004structural], and the diffusion of news feed content [bakshy2012role]. Similarly, we expect that the software-related knowledge transferred through weak ties is more valuable for the creation of innovative projects. Thus, we hypothesize that:
@@
H3. The more the informational diversity of developers’ past interactions can be attributed to weak ties, the more innovative their projects are.
III Methods
To test our hypotheses, we reconstruct three historical project-to-project interaction networks (encoding possible knowledge flows across ties of varying strength) for a large sample of open-source Python projects hosted on GitHub, compute vector representations for all the nodes, and compute measures of tie strength and tie diversity based on these networks and representations. Separately, inspired by Fang et al.’s [fang2024novelty] operationalization of software innovativeness as novel recombination of software packages, we propose and compute a novel measure of project innovativeness that is better able to capture context and is also more scalable. Finally, we bring the two parts together, using regression modeling on this longitudinal data to test our hypotheses about factors associated with higher levels of innovativeness.
III-A Initial Data Collection
For our study, we augment existing data from the World of Code (WoC) dataset [ma2019world] with additional data mined using the GitHub API. WoC is arguably the most comprehensive record of open-source projects and their commit activities. Conveniently, WoC also contains pre-computed historical package manager dependency information for all projects using the most popular programming language / package manager combinations—we use this information when computing project innovativeness.
To keep our analysis tractable, we focus only on projects primarily written in Python, i.e., projects where more than 50% of the source code files are Python files. We chose Python due to its immense popularity in open source [srinath2017python] and its widespread usage across various domains by developers from diverse backgrounds [saabith2019python]. While not allowing us to directly generalize our findings, the considerable variability in the Python ecosystem across many characteristics should ensure points of overlap with many other ecosystems.
Because many GitHub repositories are intended as code dumps rather than active development projects and to account for more of the intricacies of GitHub data [kalliamvakou2016depth], we further apply a few common sense filters. That is, we exclude forks and repositories with fewer than ten commits in total. In addition, we further exclude repositories created after 2022, so that we can observe enough of the projects’ development to make inferences about their innovativeness.
For this remaining set of projects we query the GitHub API for historical information on all their commits, issues, and stars (the latter two are not available in WoC); for each such event we record its author and timestamp. We use this data for our network construction, detailed in Section III-B below.
As a final preprocessing step, we used typical heuristics to identify and filter out bot accounts [dey2020detecting, wessel2018power], which would distort our networks and subsequent measures, as they tend to have lots of activity. To this end, we searched for keywords in user logins, such as -bot or -robot, and labeled those users as bots after manual review. Additionally, we manually inspected the profile descriptions of the top 100 most active user accounts in our sample, ordered by the number of commits, to ensure the removal of highly active bot accounts that may have escaped the naming convention heuristics.
III-B Network Construction
@@
We theorized in Section II above that 1) the interactions developers have with each other and with each other’s artifacts, i.e., the ties they form, create opportunities for knowledge transfer and facilitate social learning; and 2) “strong” and “weak” ties may play different roles and may have different effects. To operationalize these concepts, we first construct three interaction networks where every node is a project (i.e., GitHub repository), and the links between two nodes encode different interactions the developers of one project had with the other project; links are directed. We operationalize “developers” and “interactions” as follows.
III-B1 Developers—Core Developers
We consider only the actions of core developers in a focal project, identified as those contributors who authored at least five percent of all commits in the repository, and contributed a minimum of ten commits to the project. Conceptually, in open source the core developers wield the most influence over a project’s technical design decisions and development trajectory. Thus, we expect that the knowledge they (as opposed to more peripheral contributors) had access to prior to developing a focal project has the most influence on the design of that project. Pragmatically, focusing on core developers has the added advantage of keeping the analysis tractable, as popular projects may have tens of thousands of peripheral contributors.
Note that we use the benefit of hindsight here: i.e., essentially we label developers as “core” already at the start of a project, based on the amount of contributions they eventually made. Other alternatives [robillard2024communicating] that don’t rely on hindsight are possible, like considering the founder of a project. However, these have other limitations; e.g., there can be many influential developers that join a project later. Network-based operationalizations of “core” versus “periphery” have also been proposed [joblin2017classifying]. However, prior work [joblin2017classifying] demonstrated that count-based and network-based measures generally agree.
III-B2 Interactions—Commits, Issues, Stars
@@
We consider three types of actions that one can make in a project: contributing changes to the codebase by making commits, posting or participating in an issue discussion, and starring a repository. Besides varying in effort (with commits typically requiring the most effort and star gazing the least of the three), conceptually, the three should also reflect varying levels of knowledge flow from the project to the author of the action. While commits can vary in size [alali2008s], a successful commit often requires a significant understanding of the project’s technical details [von2003community], indicating a path for considerable knowledge flow between the project and the commit author. Issues threads contain discussions on various project-related topics, like feature suggestions, bug reports, and user problems [merten2015requirements]. Engaging in issue discussions can reflect a deep understanding of the project, which can be further enhanced through feedback and comments on the same thread. However, on average, we expect that participating in issue discussions requires (and reflects) a less deep understanding of the project compared to making changes to its codebase. Finally, starring a project is often done as a sign of appreciation, a bookmarking attempt, or an intent for later use [borges2018s]. Starring a repository indicates at least some awareness of the project, but on average probably much more basic understanding than the other two.

### v2

# The Strength of Weak Ties Between Open-Source Developers

@@
Abstract.
@@
In a real-world social network, weak ties (reflecting low-intensity, infrequent interactions) act as bridges and connect people to different social circles, giving them access to diverse information and opportunities that are not available within one’s immediate, close-knit vicinity. Weak ties can be crucial for creativity and innovation, as they introduce ideas and approaches that people can then combine in novel ways, leading to innovative solutions. Do weak ties facilitate creativity in software in similar ways? This paper suggests that the answer is “yes.” Concretely, we study the correlation between developers’ knowledge acquisition through three distinct interaction networks on GitHub and the innovativeness of the projects they develop, across over 37,000 Python projects hosted on GitHub. Our findings suggest that the topical diversity of projects in which developers engage, rather than the volume, correlates positively with the innovativeness of their future code. Notably, exposure through weak interactions (e.g., starring) emerges as a stronger predictor of future novelty than via strong ones (e.g., committing).
1. Introduction
Think of examples of big software innovations. You might name the Netscape browser (created the visual web and fueled mass internet adoption), the Git version control system (revolutionized collaborative software development), Hadoop MapReduce (enabled “big data” processing), Photoshop (made digital image manipulation mainstream), Netflix (killed video rental stores), Uber (“revolutionized” transportation), or perhaps more than anything else, the World Wide Web (the global information network that connected humanity and fundamentally changed how we access knowledge).
Besides their undeniable impact, what all these innovations have in common is, perhaps surprisingly, that none was highly novel in a pure sense. Instead, they succeeded by recognizing latent potential in existing components that others had overlooked or dismissed as insufficient. This reveals a fundamental paradox in software innovation: breakthroughs can emerge not only from inventing entirely new concepts, but from developing the insight to see how familiar pieces can fit together in unexpected ways.
As he remarkably candidly notes in his book “Weaving the Web” (Berners-Lee, 1999), Berners-Lee didn’t invent hypertext, markup languages, or networking protocols, but he saw how HTTP, HTML, and URLs could create a self-reinforcing web that previous hypertext systems had failed to achieve. The web browsers didn’t invent hypertext or networking either, but they made the connection between them obvious in retrospect. Git didn’t create new concepts around file differencing, cryptographic hashing, or distributed systems, but made distributed version control finally practical. Photoshop didn’t invent image processing algorithms or digital manipulation, but unified traditional darkroom techniques with mathematical transforms in an intuitive interface. Uber didn’t invent transportation technology, but cleverly orchestrated GPS, mobile payments, and dispatch algorithms that had existed separately for years. Finally, Netflix didn’t invent video streaming or recommendation systems – they combined content delivery networks with collaborative filtering at a moment when bandwidth made it viable.
What distinguishes many impactful software innovations, then, isn’t pure novelty but rather a form of combinational creativity (Boden, 2004) – the ability to perceive new relationships between existing elements and to pursue combinations that may seem counterintuitive. Many innovators aren’t necessarily the first to create the individual components, but they are the first to see the gestalt that emerges from their combination. This suggests that thinking of breakthrough innovations as pure inventions may actually mislead us about where a lot of transformational software comes from. Perhaps a highly valuable skill isn’t creating something from nothing, but developing the pattern recognition to spot which existing pieces, when properly orchestrated, can produce emergent behaviors that transcend their individual capabilities. Much real innovation, thus, lies not in the novelty of the parts, but in the non-obvious wisdom of their assembly. In fact, this process of innovation through novel recombination is not specific to software – it appears to be a fundamental pattern in virtually all domains of human innovation, and it has been theorized and studied empirically in many contexts, including business, medicine, science, and technology (Schumpeter, 1939; Uzzi et al., 2013; Hofstra et al., 2020; Shi and Evans, 2023).
But while it’s clear that innovation drives the software industry (Edison et al., 2013), and that great software engineers (among many other attributes) should be “[able to] generate novel and innovative solutions based on the context and its limitations” (Li et al., 2020), it’s still unclear how to develop one’s ability to combine existing things in clever ways. Such combinational creativity “typically requires a very rich store of knowledge, of many different kinds, and the ability to form links of many different types” (Boden, 2004). How to develop these? Although empirical evidence is scarce (Mohanani et al., 2017; Groeneveld et al., 2021), one might theorize that mixing knowledge depth with knowledge breadth is the key (da Silva et al., 2015; Turner et al., 2002), since these characteristics make software professionals more successful in general (Sonnentag, 1995; Dieste et al., 2017). According to this thinking, deep engagement with the software development process, ideally through “deliberate practice” (Baltes and Diehl, 2018), should help build expertise, while broad exploration in many different contexts should help build vocabulary.
Our paper goes one step further in this line of thinking, revealing a stronger and more nuanced effect of breadth than one might expect. Across over 37,000 Python projects on GitHub, while we confirm an association between knowledge breadth and software innovation, we find that exposure through minimal-effort interactions like starring repositories emerges as a stronger predictor of future innovation than high-effort engagements like writing code. This counterintuitive result suggests that casual, passive observation of diverse projects – essentially digital lurking – may be more valuable for sparking creative recombination than direct, intensive engagement and active learning.
Much like the unexpected advantage of weak ties in job searches, which challenged conventional wisdom of social networks at the time (Granovetter, 1973), this “strength of weak ties on GitHub” represents a paradox worth exploring: Sometimes, developers’ least-effort interactions with projects may contribute the most to their innovation capacity. Our research investigates this surprising mechanism, examining how the structural characteristics of developers’ interaction networks correlate with the innovativeness of the software artifacts they subsequently create.
Below, we draw from established theory to formulate our hypotheses (§2); propose a novel software innovativeness measure, reconstruct strong- and weak-tie interaction networks (via commits, issues, and starring) for a project’s core team, and estimate the amount and diversity of knowledge accessible to the core developers via these networks (§3); build regression models to test the association between network structural properties and software innovativeness (§4); and discuss the implications of our results (§5).
2. Theory and Hypotheses
In a nutshell, developers on the GitHub platform engage in a wide range of interactions, corresponding to ties of varying strength in a network sense, that create opportunities for knowledge acquisition via social learning. This knowledge can influence the innovativeness of the solutions they create. Before diving into the details of the mechanism, let us clarify some key constructs.
Innovation as Novel Recombination. As discussed above, we view software innovation as often emerging from novel recombinations of existing components, libraries, and patterns, following the long-standing view that “[business] innovation combines factors in a new way, or that it consists in carrying out new combinations” (Schumpeter, 1939), which has fueled much research in the social sciences (e.g., (Uzzi et al., 2013; Shi and Evans, 2023; Hofstra et al., 2020)). As a result, the line between novelty and innovation becomes fuzzy, thus we will use the two terms interchangeably, in contrast to some prior work that distinguishes them (Couger and Dengate, 1992; Beghelli and Jones, 2020).
@@ 2. Theory and Hypotheses
Creativity as Antecedent of Innovation. How to facilitate the emergence of innovation is unclear, although a common argument centers on enabling creativity as a precursor. Creativity and innovation are closely related but distinct constructs: creativity is the ability to generate novel ideas, while innovation transforms novel concepts into tangible outcomes. It is expected that a software team that is more creative has access to richer knowledge (e.g., knowledge about the problem and technical approaches to solving it), and is better able to combine pieces of that knowledge to create an innovative solution (Du Plessis, 2007). Of course, there are many facets of creativity, including as an individual personality trait (Kandler et al., 2016). In this paper we focus only on one. Specifically, we argue that creativity at the team level is a function of the knowledge networks of the team members and, while we cannot measure creativity directly, we can expect to see differences in the teams’ creative outputs (i.e., more novel software), that are associated with differences in the structure of those knowledge networks.
@@
Knowledge Acquisition via Social Learning. Historically, and especially with the advent of “social coding” platforms like GitHub, participation in open source has been rife with opportunities for social learning (Crowston and Howison, 2005), i.e., the process by which individuals acquire new knowledge, behaviors, skills, or attitudes through observation, imitation, and interaction with others. Learning, including by observing what others are doing (Yang et al., 2021), has been (Lakhani and Wolf, 2005; Krishnamurthy, 2006) and remains (Gerosa et al., 2021) among the most important motivations for people to contribute to open source. In addition, modern code hosting platforms offer many opportunities for users to interact (e.g., collaboration on a shared codebase, issue discussions, code review). Through social media-like functionality (e.g., “following” (Blincoe et al., 2016)) and many available signals that offer a high degree of transparency (Dabbish et al., 2012) (e.g., repository badges (Trockman et al., 2018)), the platforms also facilitate users quietly observing and being influenced by others’ behaviors.
@@
These interactions with other individuals (Von Krogh et al., 2003) and with the artifacts they create (Cervetti et al., 2009) serve as channels for information dissemination and knowledge exchange. Indeed, prior studies have documented many such social learning effects, including choosing an open-source license (Peng et al., 2013; Singh and Phelps, 2013), discovering and adopting emerging tools (Singer et al., 2014; Lamba et al., 2020), and learning new software design principles and programming skills by reviewing code authored by others (Nandigam et al., 2008; Alami et al., 2019).
@@
Still, if innovation requires an increased “vocabulary” of knowledge bits accessible for recombination, even high volumes of interaction might not be sufficient if the corresponding information is redundant. This can happen, e.g., when one works primarily within a narrow domain. Instead, we expect that accessing diverse knowledge acts as a catalyst for innovation, enabling individuals to integrate disparate concepts and to develop unconventional and novel products (Schumpeter and Swedberg, 2021). Outside of software engineering, Tortoriello et al. (2015) observed a positive association between employees’ access to diverse knowledge within the research departments of high-tech companies and their innovation levels. Similarly, Abdul Basit and Medase (2019) found that the integration of knowledge from both internal research teams and customers in the public sector enhances innovation. Thus, we hypothesize that:
@@
The Strength of Weak Ties. In a social network, ties can have varying strength, reflective of real-world factors like the duration of shared interactions, emotional depth, level of intimacy, and amount of reciprocal exchanges (Granovetter, 1973). In that sense, “weak” ties involve infrequent and less intimate interactions, while “strong” ties are characterized by higher frequency of interaction and intimacy.
In the late 1960s, sociologist Mark Granovetter uncovered a counterintuitive finding that would revolutionize our understanding of social networks: When searching for jobs, people relied more on casual acquaintances (weak ties) than close friends or family (strong ties). His groundbreaking theory on the “strength of weak ties” (Granovetter, 1973) revealed that these seemingly superficial connections often serve as critical bridges between different social circles, providing access to novel information unavailable within one’s immediate network. This pattern arises because individuals within the same network neighborhood have numerous opportunities for interaction and often share many mutual ties. This shared social context leads to increased similarity in behaviors, interests, and other characteristics. Consequently, the information disseminated via strong ties risks becoming less novel and less valuable (the discussion of echo chambers on social media (Kitchens et al., 2020) is a prime example of this degradation of information quality). In contrast, information acquired via weak ties, or cross-group connections, tends to originate more from individuals with diverse backgrounds and knowledge bases, making it more likely to be novel and of greater value (Granovetter, 1973). This mechanism has been extensively validated in social contexts, including generation of innovative ideas (Burt, 2004), the diffusion of news feed content (Bakshy et al., 2012), and job seeking in the digital age (Rajkumar et al., 2022).
@@
Similarly, we expect that the software-related knowledge transferred through weak ties is more valuable for the creation of innovative software projects. Thus, we hypothesize that:
@@
H3. The more the informational diversity of developers’ past interactions is due to weak ties, the more innovative their projects are.
3. Methods
Next we give a high-level overview of our approach, before diving into operationalization details.
3.1. Overview / Intuition
@@
Network Construction. We theorized above that (1) the interactions developers have with each other and with each other’s artifacts, i.e., the ties they form, create opportunities for knowledge transfer and facilitate social learning; and (2) “strong” and “weak” ties may play different roles in this knowledge diffusion process. To operationalize these concepts, we first construct three interaction networks where every node is a project (i.e., GitHub repository), and the links between two nodes encode different interactions the developers of one project had with the other project.
There are many ways in which two open-source developers can interact, both in and outside the GitHub platform. Clearly, it’s not possible to capture all interactions at scale, as projects may use, e.g., a diversity of communication channels, including private ones. Instead, we consider three representative examples of interactions that (a) vary substantially in effort, thus can be expected to reflect varying levels of knowledge flow from the target project to the author of the action (i.e., to encode person–to–project ties of varying strength, which we later project to construct our project–to–project networks), (b) are core features of the GitHub platform, thus are commonly used: making commits to a codebase (direct pushes or merged pull requests), posting or participating in issue discussions, and starring repositories.
@@
Among the three, commits typically require the most effort and starring the least. This ranking should correspond to the depth of the knowledge that may transfer as a result of each action. Intuitively, while commits vary in size and content (Alali et al., 2008), an “average” commit should require a significant understanding of the project’s technical details (Von Krogh et al., 2003), indicating a path for considerable knowledge flow between the project and the commit author. Issue threads primarily discuss feature suggestions, bug reports, and user support (Merten et al., 2015). Thus, while engaging in issue discussions can reflect a deep understanding of the project, on average we expect that it requires (and reflects) a less deep understanding of the project compared to making changes to its codebase. Finally, starring a project is often done as a sign of appreciation, a bookmarking attempt, or an intent for later use (Borges and Valente, 2018). Starring a repository indicates at least some awareness of the project, but on average probably much more basic understanding than the other two.
Note that we consider only the actions of core developers in a focal project, identified heuristically as those contributors who authored at least five percent of all commits, with a minimum of 10 commits in total. Our operationzalization is validated with alternative threshold of core developer identification in the appendix. Conceptually, in open source, core developers wield the most influence over a project’s technical decisions and development trajectory (Robles et al., 2009). Empirically, in our sample core developers also authored the vast majority (close to 90 percent) of commits importing new packages into projects (which we use to compute innovativeness, as described below). Thus, we expect that the knowledge they had access to prior to working on a focal project (as opposed to peripheral contributors) has the most influence on the design and innovativeness of that project.
As possible alternatives (Robillard et al., 2024), we considered reflecting the past interactions of only the founder of a project (which would have missed influential people joining a project later), as well as network centrality-based operationalizations of “core” versus “periphery” status (which are more precise but computationally more complex and, on average, correlate highly with count-based measures like ours (Joblin et al., 2017)). Finally, we considered capturing the past interactions of all project contributors. However, at the scale of our study (over 37,000 projects) this was infeasible considering that popular projects may have tens of thousands of peripheral contributors.
Network Measures. Using our network data, we compute variables related to the network position of each project to represent the amount and diversity of knowledge that may be accessible to its core developers through each network. For example, consider the commit-based project–to–project network. A node (project) in this network may have many outgoing connections, indicating that its core developers have collectively committed to many other projects prior to joining the focal project. That is, they have first-hand experience with the technologies used in those other projects, and they can probably draw from that experience (knowledge) now to influence the development of the focal project. Similarly, we can reason about the other two networks constructed from issue discussions and starring repositories.
How to measure the diversity of knowledge is less obvious. Conceptually, the amount and diversity of knowledge are related yet distinct. For example, a project whose core developers are connected to a large number of other projects within the network may have access to diverse knowledge. However, the actual diversity of this knowledge remains contingent on the nature of these connections – if all linked projects contain similar knowledge, diversity may be limited; e.g., if they’re all related to visualization, that doesn’t say much about experience with tensorflow.
To capture this diversity, we adopt an approach analogous to word embeddings, wherein semantic similarity between words is inferred from their vector representations (Mikolov et al., 2013b). The key ideas is that we compute project graph embeddings and use embedding distance to quantify the similarity of gainable knowledge among projects. The intuition is that nodes that frequently appear together in the same random walk while learning the embeddings represent projects that are often contributed to or interacted with by the same developers, i.e., are related in some informational sense. This could be because they share similar functionality, are used in similar contexts, or are part of the same development ecosystem.
Dimensionality Reduction. Thus far we have been treating our three networks (commits, issues, and stars) as separate. However, one can expect that the variables we compute based on these networks are correlated to some extent. Moreover, one can imagine considering other types of interactions (i.e., networks) besides our three, which would result in yet more variables. To prevent issues of multicollinearity in our subsequent regressions, which can both complicate the interpretation of the estimated coefficients and reduce statistical power (Daoud, 2017), we use Principal Component Analysis (PCA) (Wold et al., 1987) to decompose the variables representing the amount and diversity of knowledge into orthogonal components. As we show below, this produces a clear decomposition of our original variables into two components capturing strong and weak ties, which we use in our models instead of the original variables.
Project Innovativeness Measure. Fang et al. (2024) introduced a measure of a software project’s novelty, or innovativeness, as a function of the combinations of packages the project imports. For example, while certain libraries, such as numpy and tensorflow, are frequently used together, others, like numpy and requests, are less commonly combined. The measure estimates how atypical combining two packages is, and by extension how atypical is the overall set of packages being used as dependencies in a project, by comparison to what could be expected by random chance. Projects that import more atypically paired packages are deemed more innovative as they reuse packages in nontraditional ways.
We maintain this framing of innovation through novel recombination, and similarly consider packages as the unit of recombination. Packages are designed to be reused; they’re typically substantial enough to provide real functionality while remaining composable (individual functions are too small – they lack sufficient context and capability to be meaningfully recombined; entire applications are too large); and they’re usually semantically coherent, i.e., they tend to be organized around conceptual domains (e.g., requests encapsulates the entire complexity of HTTP communication in a reusable form, while pandas bundles decades of data manipulation knowledge into a coherent interface). Alternatively, one could consider something like “concepts” as the unit of recombination, as in some prior sience-of-science work (Hofstra et al., 2020). However, it wasn’t clear how to automatically extract concepts from software, so we chose packages as the unit instead.
Our measure of project innovativeness, while similar in spirit to Fang et al’s (Fang et al., 2024), is based on learning embeddings of packages and addresses two important limitations of the original measure. First, the original measure only captures pairwise combinations between packages (and aggregates up to the project level from that), neglecting the possible interactions among related packages. However, packages often form “stacks” comprised of more than two libraries that complement each other and are often reused together. Our embeddings-based approach considers more context and is expected to better capture a package’s technical use, function, and “stack.” Second, computing the original measure requires random reshuffling of the global dependency network, which is computationally prohibitive at the scale of our study. In contrast, we compute the cosine similarity of embeddings, which is faster.
Regression Analysis. Putting everything we discussed so far together, our analysis involves modeling the variation in project innovativeness across our sample as a function of the variables we discussed above, controlling for known covariates. While not causal,111Testing our hypotheses experimentally is infeasible, and identification is otherwise unclear from observational data like ours, since there is no clear “intervention.” this analysis will still allow us to test the extent to which the observational data is consistent with the causal paths theorized above (about determinants of combinatorial creativity).
In the remainder of this section we give lower-level operationalization details for the steps above, including validation checks.
3.2. Sample Selection
We combine data from World of Code (WoC) (Ma et al., 2019) and the GitHub API. WoC is arguably the most comprehensive record of open-source projects and their commit history, and it also contains historical import-based dependency information for projects in the most popular programming languages, which we use when computing project innovativeness. To keep our analysis tractable, we focus only on Python projects, i.e., the project contains more than 10 Python source code scripts. We chose Python due to its immense popularity (Srinath, 2017) and its widespread use across various domains by developers from diverse backgrounds (Saabith et al., 2019). While not allowing us to directly generalize our findings, the considerable variability in the Python ecosystem across many characteristics should ensure points of overlap with many other ecosystems.
Because many public repositories are intended as code dumps rather than active development projects, and to account for more of the intricacies of GitHub data (Kalliamvakou et al., 2016), we excluded forks and repositories with fewer than ten commits in total. In addition, we excluded repositories created before 2008 (pre GitHub) and after 2022 (which allows for at least one year of history in WoC, which we need to make inferences about project innovativeness).
For this remaining set of projects we queried the GitHub API for historical information on all their commits, issues, and stars (the latter two are not available in WoC); for each such event we recorded its author and timestamp. We used this data for our network construction, detailed in Section 3.3 below.
As a final preprocessing step, we heuristically identified and filtered out bot accounts (Dey et al., 2020; Wessel et al., 2018), which would distort our networks and subsequent measures, as they tend to have a lot of activity. To this end, we searched for keywords in user logins, such as -bot or -robot, and labeled those users as bots after manual review. Additionally, we manually inspected the profile descriptions of the top 100 most active user accounts in our sample, ordered by the number of commits, to ensure the removal of highly active bot accounts that may have escaped the naming convention heuristics.
3.3. Network Construction